OpenStudy (bmason0610):

can someone please help me? how do i solve x^2-121=0

OpenStudy (phi):

there are (at least) two ways one way is first add121 to both sides, like this \[ x^2 -121 +121= 0+121\] now simplify both sides what do you get ?

OpenStudy (bmason0610):

x^2=121.

OpenStudy (bmason0610):

but, i dont know how to simply x^2

OpenStudy (phi):

by definition \[ \sqrt{x^2} = x\] square root "undoes" the square but to keep things equal, we have to square both sides

OpenStudy (phi):

*square root both sides

OpenStudy (bmason0610):

ohh, okay, so x=11?

OpenStudy (phi):

yes, but I should mention that there are two answers to a square root people would write \[ x = \pm \sqrt{121}= \pm 11 \] which is short for x=11 or x=-11 (notice -11*-11 is +121)

OpenStudy (bmason0610):

so when you square root a problem to simplify, it's both?

OpenStudy (phi):

yes, when solving these type problems btw, in theory we would do \[ \pm\sqrt{x^2}= \pm \sqrt{121} \\ \pm x = \pm 11\] but it turns out that we only have to +/- one side to get all of the answers so people usually do the right side.

OpenStudy (bmason0610):

oh okay, that makes sense, thank you.